# A Review of Online Diffusion Policy RL Algorithms for Scalable Robotic Control

### v1

@@ 3.1.1 Model-free Online RL with DIffusion POlicy (DIPO)
3.1.2 Deep Diffusion Policy Gradient (DDiffPG)
DDiffPG [DDiffPG] extends the methodology of DIPO to actively promote multimodal behaviors. Specifically, this approach collects multiple trajectories through rollouts and extracts $M$ representative modes using hierarchical clustering [hierarchicalclustering]. Each clustered trajectory set, denoted as $\tau_{m}$, then undergoes a process of action improvement similar to Eq. 10, utilizing a mode-specific parameterized Q-function $Q_{\phi_{m}}(s,a)$ as follows:
@@ 3.1.2 Deep Diffusion Policy Gradient (DDiffPG)
3.1.3 Q-Score Matching (QSM)
Compared to DIPO and DDiffPG, QSM [qsm] proposes a more direct approach to training diffusion policies using action gradients through Q-score matching. QSM theoretically proves that aligning the score function of score-based generative models with action gradients leads to global improvement of the policy’s action-value function. Given score estimator $\mathbf{s}_{\theta}(\cdot,\cdot)$ and action gradient $\nabla_{a}Q(s,a)$, the empirical loss is:
@@ 3.2 $Q$-weighting Methods
3.2.1 Q-weighted Variational Policy Optimization (QVPO)
QVPO [qvpoerror] is the pioneering work in $Q$-weighting methods, theoretically proving that when all state-action values $Q(s,a)$ are positive for any state-action pair $(s,a)$, the Q-weighted variational bound objective of the diffusion model [ddpm] provides a tight lower bound of the policy objective in online RL. However, since $Q(s,a)$ is not always positive in general settings, QVPO proposes a weight transformation function utilizing the advantage term:
@@ 3.2.1 Q-weighted Variational Policy Optimization (QVPO)
3.2.2 Diffusion Policy Mirror Descent & Soft Diffusion Actor-Critic (DPMD & SDAC)
Similar to QVPO, Ma et al. [dpmdsdac] introduce reweighted score matching methods, which constitute a general loss family as follows:
@@ 3.2.2 Diffusion Policy Mirror Descent & Soft Diffusion Actor-Critic (DPMD & SDAC)
where $p_{k}(a^{k}|s)$ is the noise-perturbed policy with a Gaussian kernel. Note that we can compute $p_{k}(a^{k}|s)$ in a tractable space (e.g., Gaussian space) through several mathematical foundations such as Tweedie’s identity [tweedie].
Specifically, they introduce two different $Q$-weighting approaches named DPMD and SDAC, based on mirror descent policy [md] and soft actor-critic [sac], respectively. Both algorithms employ the following exponential reweighting function:
@@
3.3 Proximity-Based Methods
Several proximity-based methods (e.g., TRPO [trpo], PPO [ppo]) require direct computation of the log-likelihood of the policy $\pi(a|s)$ as follows:
@@ 3.3 Proximity-Based Methods
3.3.1 Generative Policy Optimization (GenPO)
GenPO [genpo] computes $\pi_{\theta}(a|s)$ through exact diffusion inversion [edict] combined with a doubled dummy action mechanism. Given an original action $a\in\mathbb{R}^{d}$, GenPO defines a dummy action $\tilde{a}=(x,y)\in\mathbb{R}^{2d}$ where the final action is computed as $a=\frac{x+y}{2}$. By alternately updating the two components of the dummy action, GenPO constructs an invertible flow process. This enables exact probability density computation via the change of variables theorem from normalizing flows [nice, realnvp]:
@@ 3.3.1 Generative Policy Optimization (GenPO)
3.3.2 Flow Policy Optimization (FPO)
The key innovation of FPO [fpo] is that it reformulates the policy optimization objective by replacing the intractable log-likelihood ratio $r_{\theta}$ with conditional flow matching losses $\mathcal{L}_{\text{CFM}}$. Specifically, FPO treats the flow matching objective as a surrogate for the log-likelihood in policy gradient methods. Given the conditional flow matching loss $\mathcal{L}_{\text{CFM}}$ and advantage $A(s,a)$, FPO optimizes:
@@ 3.3.2 Flow Policy Optimization (FPO)
3.4 BackPropagation Through Time (BPTT)-based Methods
These methods do not rely on the surrogate objectives typically used in diffusion models, but instead train through the entire diffusion process in a manner similar to BPTT of recurrent neural network training. To mitigate the high computational cost of BPTT, they employ strategies such as using a small number of timesteps [dime, dacer, dacerv2], or leveraging approximations through few-step generation techniques like consistency models [cpql]. However, these approaches face potential scalability issues: when extending the number of sampling steps to improve performance, they either incur increased computational costs or suffer from degraded action quality due to approximation errors.
3.4.1 Diffusion Actor-Critic with Entropy Regulator (DACER)
@@
Similar to MaxEnt RL frameworks [sql, sac], DACER introduces a mechanism to regulate the exploitation-exploration trade-off by incorporating an estimated entropy $\hat{\mathcal{H}}=\mathbb{E}_{a\sim\pi}[-\beta\log\pi(a|s)]$. To circumvent the intractability of calculating entropy directly from diffusion models, the authors leverage a tractable Gaussian Mixture Model (GMM) as a proxy. This approximated entropy $\hat{\mathcal{H}}$ is then used to update the exploration parameter $\alpha$ according to the following formulation:
@@
3.4.2 Enhanced DACER (DACERv2)
DACERv2 [dacerv2] is specifically designed to overcome the scalability bottlenecks associated with the BPTT-based approach of DACER, aiming to maintain high performance even with a significantly reduced number of diffusion steps. To achieve this, it incorporates action gradient guidance—similar to the QSM [qsm] framework—as an auxiliary objective, thereby placing DACERv2 within the action-gradient family of algorithms. Specifically, they extend the standard action gradient $\nabla_{a}Q(s,a)$ into a Q-gradient field $\nabla_{a^{k}}Q(s,a^{k})$, enabling the utilization of action gradients across all diffusion steps $k$. By leveraging this proposed Q-gradient field, the method facilitates a more sophisticated score-matching process with a time-dependent score model, analogous to QSM [qsm], thereby enhancing overall performance as follows:
@@ 3.4.2 Enhanced DACER (DACERv2)
3.4.3 Diffusion-Based Maximum Entropy RL (DIME)
Unlike DACER and DACERv2, which employ entropy terms primarily as regularizers, DIME [dime] proposes a more direct framework for training diffusion policies within the MaxEnt RL paradigm [sql]. To overcome the intractability of calculating entropy in diffusion models, this method derives a tractable lower bound for the MaxEnt objective by leveraging approximate inference techniques [berner2022optimal]. For policy improvement, DIME formulates and solves the following approximate inference problem:
@@ 3.4.3 Diffusion-Based Maximum Entropy RL (DIME)
3.4.4 Consistency Policy with Q-Learning (CPQL)
CPQL [cpql] addresses the dual challenges of inference latency and inaccurate guidance in diffusion-based RL by leveraging the consistency model [cm] for one-step action generation. CPQL formulates the policy as a parameterized consistency function $f_{\theta}(\cdot)$ that directly maps noise to actions in a single step as follows:
@@ 3.4.4 Consistency Policy with Q-Learning (CPQL)
3.4.5 Distributional Soft Actor-Critic with Diffusion Policy (DSAC-D)
Departing from prior works rooted in standard or MaxEntRL, DSAC-D [dsacd] represents the first attempt to integrate diffusion networks with Distributional RL [c51, dsac, dsact], which models the probability distribution of returns $Z(s,a)$ rather than a scalar $Q(s,a)$. To achieve a more precise approximation of the return distribution in this framework, DSAC-D introduces the Diffusion Value Network (DVN), which can model multi-peaked value distributions. Similar to Eq. 7, the DVN generates return samples through the reverse diffusion process:
@@ 3.4.5 Distributional Soft Actor-Critic with Diffusion Policy (DSAC-D)
For policy improvement, DSAC-D incorporates the entropy term $\beta\log\pi(a|s)$ from MaxEnt RL into a revised distributional Bellman equation [c51] to facilitate exploration. The loss function for the diffusion policy with distributional return under the policy $Z^{\pi}(s,a)$ is formulated as follows:
@@
Note that to estimate the entropy term $\log\pi(a|s)$, which balances exploitation and exploration, DSAC-D adopts the entropy estimation technique from DACER [dacer], utilizing Gaussian Mixture Models (GMMs) to approximate policy entropy.
3.5 Other Related Diffusion-Based RL Paradigms
@@
A closely aligned and frequently explored research direction is the Offline-to-Online Fine-tuning RL approach [dppo, parl, policydecorator, resfit, resip]. These methods typically leverage the expressive power of a diffusion policy initialized via imitation learning on a large demonstration dataset, and then subject it to policy optimization in the online RL setting. Similar to the aforementioned model-free DPRL paradigms, these approaches either propose specialized online RL frameworks tailored for diffusion policies [dppo] or utilize model-agnostic training strategies—such as decoupled policy improvement—to fine-tune large-scale pre-trained models like Vision-Language-Action (VLA) and high-capacity diffusion policies [parl]. Furthermore, recognizing the non-trivial nature of applying full RL updates to the complex diffusion network, specialized methods utilize Residual Policy Learning [rpl]. These techniques freeze the bulk of the pre-trained base diffusion policy and only learn a residual action term on top of it [policydecorator, resfit, resip].
Another branch involves Model-Based Reinforcement Learning (MBRL), where diffusion models are used to learn a world model or system dynamics rather than the policy itself. These approaches focus on training a generative model to directly estimate future states, rewards, and other dynamics given a state-action pair [dwm, world4rl] to achieve “online-like” learning through synthetic rollouts. Other works use diffusion models for synthesizing limited modalities [pgd, polygrad], such as motion planning trajectories [mpd, diffuser]. While these MBRL methods achieve “online-like” learning through synthetic rollouts, their reliance on accurately modeling the environment’s transition function fundamentally distinguishes them from the model-free policy optimization focus of this paper.
3.6 Cross-Cutting Challenges and Optimization Strategies in Online DPRL
@@
3.6.1 Addressing the Intractability of Log-Likelihood and Entropy
@@
3.6.2 Enhancing Exploitation via Batch Action Sampling
Due to the high stochasticity of the diffusion process, a single sampled action may not reliably represent the policy’s optimal intent, leading to inefficient exploitation. To mitigate this, many Online DPRL frameworks adopt Batch Action Sampling, a technique popularized by Diffusion-QL [diffusionql] in the offline setting. Methods such as QVPO, SDAC, and DPMD sample multiple action candidates for a given state and utilize them to reduce variance in gradient estimation or to select the most promising action based on the critic’s evaluation, thereby stabilizing the online learning signal.
3.6.3 Mitigating Value Overestimation Bias
As with conventional DRL, Online DPRL is susceptible to value overestimation bias, which can be exacerbated by the high-dimensional and multimodal nature of diffusion action spaces. To ensure stable value landscapes, many algorithms (e.g., DIPO, QVPO, CPQL, and DACERv2) incorporate Double Q-learning architectures [doubleql]. Furthermore, some recent approaches seek to capture the full aleatoric uncertainty of the environment by integrating Distributional Value Approximators. For instance, DSAC-D and DACERv2 leverage distributional RL principles to provide a more robust critic, which in turn leads to more reliable policy updates for the diffusion actor.
@@ 4 Analysis
4.1 Experimental Setup
4.1.1 Simulation Environment and Task Selection
We utilize the NVIDIA Isaac Lab benchmark [isaaclab], which is specifically designed as a large-scale robot learning framework. Isaac Lab’s environments consist of various tasks, primarily involving locomotion and manipulation. From these, we select 12 distinct Isaac Lab tasks spanning diverse robot specifications and task types, including 3 classic tasks, 6 locomotion tasks, and 3 manipulation tasks. For the locomotion tasks, the observation configuration consists of 2 tasks configured with only proprioception (internal state, such as joint angles, velocities), and 4 tasks configured with both proprioception and LiDAR exteroception (external state). We summarize the details related to task selection and state-action space in Tab. III.
@@ 4.1.1 Simulation Environment and Task Selection
PPO [ppo] DDPG [ddpg] TD3 [td3] SAC [sac] DIPO [dipoerror] QVPO [qvpoerror] DACER [dacer] GenPO [genpo]
@@
Humanoid 153.09 $\pm$ 22.16 14.24 $\pm$ 14.61 21.25 $\pm$ 5.45 31.32 $\pm$ 2.95 68.35 $\pm$ 11.71 4.91 $\pm$ 1.24 11.72 $\pm$ 19.01 23.36 $\pm$ 15.24
Franka Lift 25.88 $\pm$ 40.40 -1.38 $\pm$ 0.97 -0.16 $\pm$ 0.24 -4.80 $\pm$ 1.82 -0.06 $\pm$ 0.12 -0.23 $\pm$ 0.60 -0.17 $\pm$ 0.07 43.01 $\pm$ 63.83
Franka Open Drawer 92.03 $\pm$ 7.22 32.83 $\pm$ 4.15 17.79 $\pm$ 10.00 31.37 $\pm$ 11.21 51.46 $\pm$ 30.94 40.36 $\pm$ 14.62 6.51 $\pm$ 2.67 87.86 $\pm$ 17.28
Allegro Cube 15.37 $\pm$ 1.31 11.92 $\pm$ 0.21 20.98 $\pm$ 0.66 13.45 $\pm$ 0.69 16.77 $\pm$ 0.36 11.69 $\pm$ 0.45 7.67 $\pm$ 2.82 9.25 $\pm$ 2.17
Anymal D Flat 3.41 $\pm$ 1.46 -1.83 $\pm$ 5.32 0.29 $\pm$ 1.79 3.80 $\pm$ 0.23 1.55 $\pm$ 1.42 -4.92 $\pm$ 6.13 -16.41 $\pm$ 22.85 18.53 $\pm$ 0.99
@@
G1 Rough 10.41 $\pm$ 1.20 -6.24 $\pm$ 1.65 -6.76 $\pm$ 0.58 -7.76 $\pm$ 0.43 -4.09 $\pm$ 0.26 -3.57 $\pm$ 0.66 -4.38 $\pm$ 2.44 -5.94 $\pm$ 1.76
H1 Rough 3.73 $\pm$ 3.60 -7.90 $\pm$ 1.49 -7.32 $\pm$ 1.21 -8.96 $\pm$ 0.46 -4.69 $\pm$ 0.27 -3.98 $\pm$ 0.01 -7.30 $\pm$ 1.86 4.17 $\pm$ 9.49
Cassie Rough 17.87 $\pm$ 3.14 2.23 $\pm$ 2.84 0.46 $\pm$ 5.45 0.32 $\pm$ 2.12 -1.69 $\pm$ 4.20 -1.10 $\pm$ 3.15 -3.77 $\pm$ 2.82 10.85 $\pm$ 1.42
Mean Rank 1.92 6.58 5.50 4.33 3.58 5.33 6.67 2.08
4.1.2 Baseline Algorithms
Our benchmarking suite consists of six representative algorithms, including both classical online RL and online DPRL approaches. We select four representative methods for our evaluation: Proximal Policy Optimization (PPO) [ppo], a proximity-based trust region on-policy method; Deep Deterministic Policy Gradient (DDPG) [ddpg], a fundamental off-policy algorithm for continuous action spaces; Twin Delayed Deep Deterministic Policy Gradient (TD3) [td3], which improves upon DDPG by addressing overestimation bias; and Soft Actor-Critic (SAC) [sac], an off-policy framework based on the maximum entropy RL principle. For online DPRL approaches, we select one representative method from each of the four categories in our taxonomy: DIPO [dipoerror] for action-gradient methods, QVPO [qvpoerror] for Q-weighting approaches, GenPO [genpo] for proximity-based methods (also incorporates BPTT), and DACER [dacer] for BPTT-based approaches.
Since manually tuning hyperparameters for the proposed method may introduce implicit bias, we unify general hyperparameters (e.g., model hidden states size, replay buffer size of off-policy algorithms, discount factor) while preserving each method’s unique or sensitive hyperparameters (e.g., number of diffusion steps, expected entropy) to the greatest extent possible. Detailed hyperparameter settings can be found in the Appendix.
4.1.3 Experimental Design

### v2

@@ III-A1 Model-free Online RL with DIffusion POlicy (DIPO)
III-A2 Deep Diffusion Policy Gradient (DDiffPG)
DDiffPG [49] extends the methodology of DIPO to actively promote multimodal behaviors. Specifically, this approach collects multiple trajectories through rollouts and extracts $M$ representative modes using hierarchical clustering [66]. Each clustered trajectory set, denoted as $\tau_{m}$, then undergoes a process of action improvement similar to Eq. 10, utilizing a mode-specific parameterized Q-function $Q_{\phi_{m}}(s,a)$ as follows:
@@ III-A2 Deep Diffusion Policy Gradient (DDiffPG)
III-A3 Q-Score Matching (QSM)
Compared to DIPO and DDiffPG, QSM [71] proposes a more direct approach to training diffusion policies using action gradients through Q-score matching. QSM theoretically proves that aligning the score function of score-based generative models with action gradients leads to global improvement of the policy’s action-value function. Given score estimator $\mathbf{s}_{\theta}(\cdot,\cdot)$ and action gradient $\nabla_{a}Q(s,a)$, the empirical loss is:
@@ III-B $Q$-weighting Methods
III-B1 Q-weighted Variational Policy Optimization (QVPO)
QVPO [21] is the pioneering work in $Q$-weighting methods, theoretically proving that when all state-action values $Q(s,a)$ are positive for any state-action pair $(s,a)$, the Q-weighted variational bound objective of the diffusion model [40] provides a tight lower bound of the policy objective in online RL. However, since $Q(s,a)$ is not always positive in general settings, QVPO proposes a weight transformation function utilizing the advantage term:
@@ III-B1 Q-weighted Variational Policy Optimization (QVPO)
III-B2 Diffusion Policy Mirror Descent & Soft Diffusion Actor-Critic (DPMD & SDAC)
Similar to QVPO, Ma et al. [57] introduce reweighted score matching methods, which constitute a general loss family as follows:
@@ III-B2 Diffusion Policy Mirror Descent & Soft Diffusion Actor-Critic (DPMD & SDAC)
where $p_{k}(a^{k}|s)$ is the noise-perturbed policy with a Gaussian kernel. Note that we can compute $p_{k}(a^{k}|s)$ in a tractable space (e.g., Gaussian space) through several mathematical foundations such as Tweedie’s identity [29].
Specifically, they introduce two different $Q$-weighting approaches named DPMD and SDAC, based on mirror descent policy [90] and soft actor-critic [37], respectively. Both algorithms employ the following exponential reweighting function:
@@
III-C Proximity-Based Methods
Several proximity-based methods (e.g., TRPO [75], PPO [77]) require direct computation of the log-likelihood of the policy $\pi(a|s)$ as follows:
@@ III-C Proximity-Based Methods
III-C1 Generative Policy Optimization (GenPO)
GenPO [22] computes $\pi_{\theta}(a|s)$ through exact diffusion inversion [94] combined with a doubled dummy action mechanism. Given an original action $a\in\mathbb{R}^{d}$, GenPO defines a dummy action $\tilde{a}=(x,y)\in\mathbb{R}^{2d}$ where the final action is computed as $a=\frac{x+y}{2}$. By alternately updating the two components of the dummy action, GenPO constructs an invertible flow process. This enables exact probability density computation via the change of variables theorem from normalizing flows [25, 26]:
@@ III-C1 Generative Policy Optimization (GenPO)
III-C2 Flow Policy Optimization (FPO)
The key innovation of FPO [61] is that it reformulates the policy optimization objective by replacing the intractable log-likelihood ratio $r_{\theta}$ with conditional flow matching losses $\mathcal{L}_{\text{CFM}}$. Specifically, FPO treats the flow matching objective as a surrogate for the log-likelihood in policy gradient methods. Given the conditional flow matching loss $\mathcal{L}_{\text{CFM}}$ and advantage $A(s,a)$, FPO optimizes:
@@ III-C2 Flow Policy Optimization (FPO)
III-D BackPropagation Through Time (BPTT)-based Methods
These methods do not rely on the surrogate objectives typically used in diffusion models, but instead train through the entire diffusion process in a manner similar to BPTT of recurrent neural network training. To mitigate the high computational cost of BPTT, they employ strategies such as using a small number of timesteps [16, 95, 96], or leveraging approximations through few-step generation techniques like consistency models [18]. However, these approaches face potential scalability issues: when extending the number of sampling steps to improve performance, they either incur increased computational costs or suffer from degraded action quality due to approximation errors.
III-D1 Diffusion Actor-Critic with Entropy Regulator (DACER)
@@
Similar to MaxEnt RL frameworks [36, 37], DACER introduces a mechanism to regulate the exploitation-exploration trade-off by incorporating an estimated entropy $\hat{\mathcal{H}}=\mathbb{E}_{a\sim\pi}[-\beta\log\pi(a|s)]$. To circumvent the intractability of calculating entropy directly from diffusion models, the authors leverage a tractable Gaussian Mixture Model (GMM) as a proxy. This approximated entropy $\hat{\mathcal{H}}$ is then used to update the exploration parameter $\alpha$ according to the following formulation:
@@
III-D2 Enhanced DACER (DACERv2)
DACERv2 [96] is specifically designed to overcome the scalability bottlenecks associated with the BPTT-based approach of DACER, aiming to maintain high performance even with a significantly reduced number of diffusion steps. To achieve this, it incorporates action gradient guidance—similar to the QSM [71] framework—as an auxiliary objective, thereby placing DACERv2 within the action-gradient family of algorithms. Specifically, they extend the standard action gradient $\nabla_{a}Q(s,a)$ into a Q-gradient field $\nabla_{a^{k}}Q(s,a^{k})$, enabling the utilization of action gradients across all diffusion steps $k$. By leveraging this proposed Q-gradient field, the method facilitates a more sophisticated score-matching process with a time-dependent score model, analogous to QSM [71], thereby enhancing overall performance as follows:
@@ III-D2 Enhanced DACER (DACERv2)
III-D3 Diffusion-Based Maximum Entropy RL (DIME)
Unlike DACER and DACERv2, which employ entropy terms primarily as regularizers, DIME [16] proposes a more direct framework for training diffusion policies within the MaxEnt RL paradigm [36]. To overcome the intractability of calculating entropy in diffusion models, this method derives a tractable lower bound for the MaxEnt objective by leveraging approximate inference techniques [10]. For policy improvement, DIME formulates and solves the following approximate inference problem:
@@ III-D3 Diffusion-Based Maximum Entropy RL (DIME)
III-D4 Consistency Policy with Q-Learning (CPQL)
CPQL [18] addresses the dual challenges of inference latency and inaccurate guidance in diffusion-based RL by leveraging the consistency model [81] for one-step action generation. CPQL formulates the policy as a parameterized consistency function $f_{\theta}(\cdot)$ that directly maps noise to actions in a single step as follows:
@@ III-D4 Consistency Policy with Q-Learning (CPQL)
III-D5 Distributional Soft Actor-Critic with Diffusion Policy (DSAC-D)
Departing from prior works rooted in standard or MaxEntRL, DSAC-D [54] represents the first attempt to integrate diffusion networks with Distributional RL [20, 59, 27], which models the probability distribution of returns $Z(s,a)$ rather than a scalar $Q(s,a)$. To achieve a more precise approximation of the return distribution in this framework, DSAC-D introduces the Diffusion Value Network (DVN), which can model multi-peaked value distributions. Similar to Eq. 7, the DVN generates return samples through the reverse diffusion process:
@@ III-D5 Distributional Soft Actor-Critic with Diffusion Policy (DSAC-D)
For policy improvement, DSAC-D incorporates the entropy term $\beta\log\pi(a|s)$ from MaxEnt RL into a revised distributional Bellman equation [20] to facilitate exploration. The loss function for the diffusion policy with distributional return under the policy $Z^{\pi}(s,a)$ is formulated as follows:
@@
Note that to estimate the entropy term $\log\pi(a|s)$, which balances exploitation and exploration, DSAC-D adopts the entropy estimation technique from DACER [95], utilizing Gaussian Mixture Models (GMMs) to approximate policy entropy.
III-E Other Related Diffusion-Based RL Paradigms
@@
A closely aligned and frequently explored research direction is the Offline-to-Online Fine-tuning RL approach [73, 60, 105, 6, 7]. These methods typically leverage the expressive power of a diffusion policy initialized via imitation learning on a large demonstration dataset, and then subject it to policy optimization in the online RL setting. Similar to the aforementioned model-free DPRL paradigms, these approaches either propose specialized online RL frameworks tailored for diffusion policies [73] or utilize model-agnostic training strategies—such as decoupled policy improvement—to fine-tune large-scale pre-trained models like Vision-Language-Action (VLA) and high-capacity diffusion policies [60]. Furthermore, recognizing the non-trivial nature of applying full RL updates to the complex diffusion network, specialized methods utilize Residual Policy Learning [79]. These techniques freeze the bulk of the pre-trained base diffusion policy and only learn a residual action term on top of it [105, 6, 7].
Another branch involves Model-Based Reinforcement Learning (MBRL), where diffusion models are used to learn a world model or system dynamics rather than the policy itself. These approaches focus on training a generative model to directly estimate future states, rewards, and other dynamics given a state-action pair [24, 45] to achieve “online-like” learning through synthetic rollouts. Other works use diffusion models for synthesizing limited modalities [43, 74], such as motion planning trajectories [15, 44]. While these MBRL methods achieve “online-like” learning through synthetic rollouts, their reliance on accurately modeling the environment’s transition function fundamentally distinguishes them from the model-free policy optimization focus of this paper.
III-F Cross-Cutting Challenges and Optimization Strategies in Online DPRL
@@
III-F1 Addressing the Intractability of Log-Likelihood and Entropy
@@
III-F2 Enhancing Exploitation via Batch Action Sampling
Due to the high stochasticity of the diffusion process, a single sampled action may not reliably represent the policy’s optimal intent, leading to inefficient exploitation. To mitigate this, many Online DPRL frameworks adopt Batch Action Sampling, a technique popularized by Diffusion-QL [97] in the offline setting. Methods such as QVPO, SDAC, and DPMD sample multiple action candidates for a given state and utilize them to reduce variance in gradient estimation or to select the most promising action based on the critic’s evaluation, thereby stabilizing the online learning signal.
III-F3 Mitigating Value Overestimation Bias
As with conventional DRL, Online DPRL is susceptible to value overestimation bias, which can be exacerbated by the high-dimensional and multimodal nature of diffusion action spaces. To ensure stable value landscapes, many algorithms (e.g., DIPO, QVPO, CPQL, and DACERv2) incorporate Double Q-learning architectures [92]. Furthermore, some recent approaches seek to capture the full aleatoric uncertainty of the environment by integrating Distributional Value Approximators. For instance, DSAC-D and DACERv2 leverage distributional RL principles to provide a more robust critic, which in turn leads to more reliable policy updates for the diffusion actor.
@@ IV Analysis
IV-A Experimental Setup
IV-A1 Simulation Environment and Task Selection
We utilize the NVIDIA Isaac Lab benchmark [62], which is specifically designed as a large-scale robot learning framework. Isaac Lab’s environments consist of various tasks, primarily involving locomotion and manipulation. From these, we select 12 distinct Isaac Lab tasks spanning diverse robot specifications and task types, including 3 classic tasks, 6 locomotion tasks, and 3 manipulation tasks. For the locomotion tasks, the observation configuration consists of 2 tasks configured with only proprioception (internal state, such as joint angles, velocities), and 4 tasks configured with both proprioception and LiDAR exteroception (external state). We summarize the details related to task selection and state-action space in Tab. III.
@@ IV-A1 Simulation Environment and Task Selection
PPO [77] DDPG [50] TD3 [33] SAC [37] DIPO [103] QVPO [21] DACER [95] GenPO [22]
@@
Humanoid 153.09 $\pm$ 22.16 14.24 $\pm$ 14.61 21.25 $\pm$ 5.45 31.32 $\pm$ 2.95 68.35 $\pm$ 11.71 4.91 $\pm$ 1.24 11.72 $\pm$ 19.01 127.33 $\pm$ 18.33
Franka Lift 25.88 $\pm$ 40.40 -1.38 $\pm$ 0.97 -0.16 $\pm$ 0.24 -4.80 $\pm$ 1.82 -0.06 $\pm$ 0.12 -0.23 $\pm$ 0.60 -0.17 $\pm$ 0.07 78.82 $\pm$ 71.01
Franka Open Drawer 92.03 $\pm$ 7.22 32.83 $\pm$ 4.15 17.79 $\pm$ 10.00 31.37 $\pm$ 11.21 51.46 $\pm$ 30.94 40.36 $\pm$ 14.62 6.51 $\pm$ 2.67 82.64 $\pm$ 23.22
Allegro Cube 15.37 $\pm$ 1.31 11.92 $\pm$ 0.21 20.98 $\pm$ 0.66 13.45 $\pm$ 0.69 16.77 $\pm$ 0.36 11.69 $\pm$ 0.45 7.67 $\pm$ 2.82 12.23 $\pm$ 1.66
Anymal D Flat 3.41 $\pm$ 1.46 -1.83 $\pm$ 5.32 0.29 $\pm$ 1.79 3.80 $\pm$ 0.23 1.55 $\pm$ 1.42 -4.92 $\pm$ 6.13 -16.41 $\pm$ 22.85 12.55 $\pm$ 5.07
@@
G1 Rough 10.41 $\pm$ 1.20 -6.24 $\pm$ 1.65 -6.76 $\pm$ 0.58 -7.76 $\pm$ 0.43 -4.09 $\pm$ 0.26 -3.57 $\pm$ 0.66 -4.38 $\pm$ 2.44 17.15 $\pm$ 1.24
H1 Rough 3.73 $\pm$ 3.60 -7.90 $\pm$ 1.49 -7.32 $\pm$ 1.21 -8.96 $\pm$ 0.46 -4.69 $\pm$ 0.27 -3.98 $\pm$ 0.01 -7.30 $\pm$ 1.86 16.02 $\pm$ 1.05
Cassie Rough 17.87 $\pm$ 3.14 2.23 $\pm$ 2.84 0.46 $\pm$ 5.45 0.32 $\pm$ 2.12 -1.69 $\pm$ 4.20 -1.10 $\pm$ 3.15 -3.77 $\pm$ 2.82 8.82 $\pm$ 2.45
Mean Rank 1.83 5.83 5.08 5.08 4.08 5.83 6.58 1.67
IV-A2 Baseline Algorithms
Our benchmarking suite consists of six representative algorithms, including both classical online RL and online DPRL approaches. We select four representative methods for our evaluation: Proximal Policy Optimization (PPO) [77], a proximity-based trust region on-policy method; Deep Deterministic Policy Gradient (DDPG) [50], a fundamental off-policy algorithm for continuous action spaces; Twin Delayed Deep Deterministic Policy Gradient (TD3) [33], which improves upon DDPG by addressing overestimation bias; and Soft Actor-Critic (SAC) [37], an off-policy framework based on the maximum entropy RL principle. For online DPRL approaches, we select one representative method from each of the four categories in our taxonomy: DIPO [103] for action-gradient methods, QVPO [21] for Q-weighting approaches, GenPO [22] for proximity-based methods (also incorporates BPTT), and DACER [95] for BPTT-based approaches.
Since manually tuning hyperparameters for the proposed method may introduce implicit bias, we unify general hyperparameters (e.g., model size, replay buffer size, discount factor) while preserving each method’s unique or sensitive hyperparameters (e.g., number of diffusion steps, expected entropy) to the greatest extent possible. Detailed hyperparameter settings can be found in the Appendix.
IV-A3 Experimental Design